such stuff as dreams are made of:
how much of you is but vacuum? outside of the nucleus an atom contain mostly an empty space. estimate the fraction of the volume of an atom that is occupied by the nucleus. assume that a typical atom has a linear dimension of about 0.1nm and that a typical nuclear dimension is about 1 fm (note that this problem introduces some new prefixes: n is nano or 10^-9, p is pico or 10^-12, f is femto or 10^-15)
give answer as a power of ten.
1
Expert's answer
2020-04-06T09:26:10-0400
"\frac{V_n}{V_a}=\frac{a_n^3}{a_a^3}"

"\frac{V_n}{V_a}=\frac{(10^{-15})^3}{(0.1\cdot10^{-9})^3}=10^{-15}"